# MemoFurther Readings

### v1

@@ 4. Characterization of coordinated users
"Rising Costs Highlight Challenges for American Families Cost of rent, energy, and other essentials surged in May In an alarming trend, the cost of essentials such as rent, energy, and groceries continues to surge, underscoring the persistent financial challenges faced by American families. While the overall Consumer Price Index (CPI) showed a slight stabilization, the specifics reveal a stark reality of escalating living expenses in our Constitutional Republic. Analyzing the Numbers The CPI indicated a 3.3% rise in overall inflation compared to the previous year. Although this marks a slight decrease from April’s 3.4% and a significant drop from the 9.1% peak in June 2022, it remains well above the Federal Reserve’s target rate of 2%. This persistent inflation underscores the ongoing economic strain on American households. Historical Context It’s noteworthy that during the four years of Donald Trump’s presidency, the average inflation rate was maintained at a modest 1.9%. This comparison highlights a more stable economic period and suggests a need for policies that can effectively manage inflation without compromising the financial well-being of citizens. Essential Expenses on the Rise A closer examination of the May report reveals substantial increases in essential costs: Rent: Up by 5.4% Mortgage: Up by 5.6% Hospital services: Up by 7.2% Car insurance: Up by 20.3% Electricity: Up by 5.9% Ground beef: Up by 4.9% Steak: Up by 5.7% Bacon: Up by 6.9% Hot dogs: Up by 7.3% These increases in essential goods and services strain the budgets of American families, making everyday living increasingly unaffordable. Public Sentiment A recent poll reflects the public’s discontent, with only 31% of voters approving of President Biden’s handling of inflation, while a significant 61% disapprove. This sentiment underscores the urgent need for effective economic policies that address the real concerns of the populace. Administration’s Response The Biden administration continues to assert progress in combating inflation. A statement from the White House on social media claimed: "Today’s report shows continued progress in lowering inflation. President Biden knows that costs are still too high for many families and we still have a lot more to do. That’s why he will keep fighting to lower drug costs, grocery prices, and energy bills." Critical Perspective However, many Americans find these assurances lacking. The everyday experience at grocery stores and gas stations starkly contrasts with the administration’s optimistic declarations, leading to a disconnect between government rhetoric and public reality. As a Constitutional Republic dedicated to ensuring democracy and the well-being of its citizens, it is imperative that our government implements policies that stabilize the economy and reduce the financial burden on American families. The rising costs of essential goods and services are a pressing concern that requires immediate and effective action to safeguard the economic future of our nation. [Source: Poll, Whitehouse on X, BLS] https://d3nkl3psvxxpe9.cloudfront.net/documents/econTabReport_maqVHQt.pdf https://x.com/WhiteHouse/status/1800957041390792843 https://www.bls.gov/news.release/pdf/cpi.pdf #GoRightNews Shared by Peter Boykin - American Political Commentator / Citizen Journalist / Activist / Constitutionalist for Liberty Web: https://PeterBoykin.com Kick: http://Kick.com/PeterBoykin YouTube: https://youtube.com/PeterBoykinForAmerica Twitter: https://twitter.com/GoRightNews Telegram: http://t.me/realpeterboykin Rumble: http://Rumble.com/GoRightNews Like the Content? Please Support! - Go Right News: Stripe: https://gorightnews.com/donations/support-gorightnews/ Cash App: http://cash.app/\$PeterBoykin1"
Table 14. Top-2 posts by total engagement for coordinated pages in Facebook
@@
"!You are a Golden Child! A Golden Child 94 Unicorn 94 Third Eye 94 Covenant 94 Harmony 94 Praise God 94 Divine Gene 94 Nine Six 94 Blue Eyes 94 Indigo Child 94 Ultra Maga 94 John John 94 White Hat 94 American Eagle 94 Carry On 94 New Earth 94 Pineapple 94 PassionForFruit"
@@
"#CAWildfireSituationUpdate as of 6-16-2024, 5:03 P.M. PST. #PostFire - 12,266 acres, 2% containment #HesperiaFire - 1,330, 7% #JunesFire - 1,076, 70% #JacksonFire - 876, ?#HernandezFire - 600, 25% #MaxFire - 500, 0% #LisaFire - 350, 0% #PointFire 150, 0%"
@@
"What is it like to be wiser than your Creator? Job tried that and later repented in dust and ashes. Job was a wise and blessed man of faith."
@@
"What was the true history behind all of these melted ruins? Certainly not the result of erosion, but perhaps a plasma storm? The results of a sudden flip in our polar magnetic electric field perhaps."
@@
"WAVELAND, Mississippi A boil advisory was issued Mon afternoon dt burst in the main water line. NASHVILLE, Indiana A boil advisory was issued Mon after a water main break on Honeysuckle Ln. MONTICELLO, Kentucky A boil advisory was issued Mon for the East Hwy 92 area dt a water main break. MILAN, Ohio Seminary Rd bw Perrin Rd/Broad St in Milan is closed dt a water main break. SASKATOON, SK - Canada North Park Wilson School is closed dt a water main break. NEW CASTLE CO., Delaware Shipley Rd at Foulk Rd was closed Mon night dt a water main break. MONTGOMERY CO., Maryland Dt a water main break Mon, a portion of the NW Branch Stream Valley Park near Highwood Terrace was undergoing repairs. SIOUX FALLS, South Dakota A water main break Mon afternoon caused flooding in downtown. BOURBON CO., Kansas There is a water main break at the Bourbon Co Transfer Station. FORTUNA, California Water will be shut off Wed to repair a broken water main on S Fortuna Blvd."
@@
Table 15. Top-5 posts by total engagement for cross-platform coordinated accounts
5. Topic Analysis
We report the most representative posts, tweets and messages—i.e., Facebook, Twitter/$\mathbb{X}$, Telegram—for each topic in Table 16, 17 and 18, as computed by BERTopic.
Topic Representative Posts
late-show
against-trump "Stephen Colbert Speaks Truth to Trump’s Most Damaging Lie - #stephencolbert #TheLateShow #Trump #DonaldTrump #MAGA #immigrants" "Stephen Colbert Speaks Truth to Trump’s Most Damaging Lie - Rick Strom #stephencolbert #TheLateShow #Trump #DonaldTrump #MAGA #immigrants"
barron-trump-rnc "Former President Donald Trump said his youngest son Barron Trump was 17-years-old in an interview, when in fact he is 18 and a nominate delegate for the RNC." "Former President Donald Trump’s 18-year-old son, Barron Trump, is serving as a delegate to the Republican National Convention."
2020-election-acceptance "Republican National Committee co-chair Lara Trump said it’s ’obvious’ that Donald Trump has accepted the 2020 election results, when in fact, he has not." "Republican National Committee co-chair Lara Trump said it’s ’obvious’ that Donald Trump has accepted the 2020 election results, when in fact, he has not."
trump-trial-dayoff "Judge Juan Merchan gave Donald Trump the day off from his hush money trial to attend Barron Trump’s graduation ceremony, but evidence suggests that he is attending political fundraiser." "Judge Juan Merchan gave Donald Trump the day off from his hush money trial to attend Barron Trump’s graduation ceremony, but evidence suggests that he is attending political fundraiser."
jacked-up-biden "Rep. Greg Murphy, a Republican, says he has evidence that President Joe Biden was using "something" before his popular State of the Union address." "Rep. Greg Murphy, a Republican, says he has evidence that President Joe Biden was using "something" before his popular State of the Union address."
trump-classified-documents "More bombshell evidence dropped against former President Donald Trump in his classified documents case, just as Judge Aileen Cannon delayed the trial indefinitely." "More bombshell evidence dropped against former President Donald Trump in his classified documents case, just as Judge Aileen Cannon delayed the trial indefinitely."
rfk "Exciting Announcement! Join us for a special interview with independent presidential candidate RFK Jr. on Monday, June 17, immediately following The Young Turks! Tune in on tyt.com/live YouTube, Facebook, or Twitch! TYT members get exclusive access to a members-only Bonus Episode after the interview, where Cenk will ask even more personal questions and dive deeper into the issues! Sign up as a member today at to join the discussion! Don’t miss this unique opportunity to engage with crucial political dialogue and watch TYT hold a presidential candidate accountable! See you there!" "Exciting Announcement! Join us for a special interview with independent presidential candidate RFK Jr. on Monday, June 17, immediately following The Young Turks! Tune in on tyt.com/live YouTube, Facebook, or Twitch! TYT members get exclusive access to a members-only Bonus Episode after the interview, where Cenk will ask even more personal questions and dive deeper into the issues! Sign up as a member today at to join the discussion! Don’t miss this unique opportunity to engage with crucial political dialogue and watch TYT hold a presidential candidate accountable! See you there!"
coulter-ramaswamy "Conservative Ann Coulter told former Republican presidential candidate Vivek Ramaswamy said while she agrees with him on many issues, she wouldn’t vote for him because he’s “an Indian.”" "Conservative Ann Coulter told former Republican presidential candidate Vivek Ramaswamy said while she agrees with him on many issues, she wouldn’t vote for him because he’s “an Indian.”"
trump-falsifying-records "Fox News’ Jeanine Pirro and conservative commentator Charlie Kirk are melting down over former President Donald Trump’s 34-count conviction for falsifying business records in his hush money trial."
bannon-sentence "Donald Trump adviser Steve Bannon proclaimed that "́no one" will shut him up as he was ordered to report to prison for refusing to testify before congress."
Table 16. Topics and their representative posts on Facebook
@@ 5. Topic Analysis
rfk "If Kennedy iselected it will cost the evil doers over a trillion dollars in the first year. If Kennedy is in the debate it will show Biden as sinile, Trump as weak and Bobby will be elected. So they won’t let Bobby in but will compensate CNN and anyone that might go to jail." "I’ve never been more convinced that both Trump and Biden fear Kennedy. The uniparty has no soul, and they hate democracy. We will never forget their attempts to silence us. Kennedy FTW"
@@
debate "Biden’s really not up to debate is what this means. Whenever a question is asked, Biden’s promoters will either tell the answer in an earplug or use a prompter screen on the podium so Biden can read the answer. This is funny." "Trump has guts though. They tried to make it as unfriendly for him as possible, and he still is going to do it. I’m sure they were hoping he would say no way, and then Biden could brag about Trump being afraid to debate him."
@@
covid-vaccine "Remember this on election day, remember All the "suddenly died" family and friends! Biden did this, he made vaccines mandatory! Just one of many bad Biden decisions!" "Trump was during Pandemic; FAILED to protect the American people when he knew damn well how infectious; deadly Covid Virus was, instead Trump told America "Covid was a Hoax"; failed to create a National Suppression Plan to save lives 700KAmericans DIED under Trump!"
@@
bowman-latimer "Bowman called Biden a liar. Bowman voted against the Infrastructure Act. Bowman even voted against the Debt Limit increase that could have jeopardized Social Security. Bowman did not earn another term as a Democrat." "Bowman is his worst enemy. His and Rape denying has turned his district against him. His Votes AGAINST the Biden Administration’s Progressive policies are NOT helping his district! Vote !"
@@
border-security "with your lies. It is that BLOCKED Bipartisan Border Security Bill bc told GOP to "Kill The Bill". publicly stated it. The Border Crisis is NOW all on for Blocking Border Bill to SECURE THE BORDER!" "This bill was a sham and would not close border,but would allow more to take American jobs!! Donalds reiterates why GOP rejected ‘bipartisan’ border bill to head off potential debate talking point"
@@
epstein-files "His flights with Epstein were with family present and years before he had an island. He’s also the only one saying anything about releasing the logs, Biden sure isn’t, and Trump took umbrage when DeSantis started advocating for their release for some strange reason…" " just another day in ’Murica, home of the child sex slaves. Trump made these folks feel they deserved Matt Gaetz’ job, Trumps’ job or Epstein’s job if they just ’networked’ enough. Investigate the real ’satanic’ sex cult that police and politicians hide."
@@
student-loan "This administration and president Biden seem to be answerable to nobody between giving money to student loan forgiveness even after the court ruled against them not getting Kennedy secret service protectiin, the list goes on no accountability." ""Summary: We estimate that President Bidenś recently announced "New Plans" to provide relief to student borrowers will cost \$84 billion, in addition to the \$475 billion that we previously estimated for President Bidenś SAVE plan.""
@@
hunter-biden "Hunter Biden’s laptop has nothing to do with the false Electoral College votes sent in for the 2020 election. Also, Alexander Smirnov (an FBI informant) was INDICTED for his false accusations about Barisma and the Bidens." "Actions speak louder than how they phrased something – especially when the actions are repeated over and over again. Their lies are proven in this week’s Hunter Biden trial where the laptop is used as evidence."
@@
blacks-against-biden "People aren’t paying attention and don’t expect him to be the President and don’t know about Biden’s accomplishments. That’s what the polls show now but that’s gonna change. Trump isn’t gonna win 20% of the black vote. No were here near it" "’At this stage in his presidency (3.5 years in) Biden must stop trying to convince Black voters his policies have made the US a significantly better place. Instead he should spend everything he’s got highlighting how anti-Black Trump is and how quickly any past gains will be lost."
Table 17. Topics and their representative posts on Twitter
@@
vaccine-skepticism "BREAKING: US doctors and scientists are currently investigating whether the COVID-19 virus is to blame for an “unusual pattern” of rare and deadly cancers that have been popping up in the wake of the pandemic.GeneralMCNews" "You, the unvaccinated are a special group of people. You stood your ground. You did not let the pressures of society sway you. You withstood the harshest discrimination seen in modern times. You were excluded from society. Some of you lost friendships and relationships with family members. You are the best that society has to offer. You used your analytical judgment and common sense. You are the tree that withstood the hurricane. You are superheroes! CovidVaccineTruth CovidVaccineTruthChat"
@@
flat-earth-conspiracy "The Encyclopedia of Freemasonry reveals that the Earth is indeed flat. Organizations like The Freemasons know the truth about God and our flat earth but they choose to keep it hidden." "Flat Earth was the cosmology of our ancestors who had direct contact with the Divine because they were closer to the Divine than we are now. That is how the ancient’s knew things despite not having access to technology we have now."
@@
Table 18. Topics and their representative posts on Telegram

### v2

@@ Appendix D Characterization of coordinated users
”Rising Costs Highlight Challenges for American Families Cost of rent, energy, and other essentials surged in May In an alarming trend, the cost of essentials such as rent, energy, and groceries continues to surge, underscoring the persistent financial challenges faced by American families. While the overall Consumer Price Index (CPI) showed a slight stabilization, the specifics reveal a stark reality of escalating living expenses in our Constitutional Republic. Analyzing the Numbers The CPI indicated a 3.3% rise in overall inflation compared to the previous year. Although this marks a slight decrease from April’s 3.4% and a significant drop from the 9.1% peak in June 2022, it remains well above the Federal Reserve’s target rate of 2%. This persistent inflation underscores the ongoing economic strain on American households. Historical Context It’s noteworthy that during the four years of Donald Trump’s presidency, the average inflation rate was maintained at a modest 1.9%. This comparison highlights a more stable economic period and suggests a need for policies that can effectively manage inflation without compromising the financial well-being of citizens. Essential Expenses on the Rise A closer examination of the May report reveals substantial increases in essential costs: Rent: Up by 5.4% Mortgage: Up by 5.6% Hospital services: Up by 7.2% Car insurance: Up by 20.3% Electricity: Up by 5.9% Ground beef: Up by 4.9% Steak: Up by 5.7% Bacon: Up by 6.9% Hot dogs: Up by 7.3% These increases in essential goods and services strain the budgets of American families, making everyday living increasingly unaffordable. Public Sentiment A recent poll reflects the public’s discontent, with only 31% of voters approving of President Biden’s handling of inflation, while a significant 61% disapprove. This sentiment underscores the urgent need for effective economic policies that address the real concerns of the populace. Administration’s Response The Biden administration continues to assert progress in combating inflation. A statement from the White House on social media claimed: ”Today’s report shows continued progress in lowering inflation. President Biden knows that costs are still too high for many families and we still have a lot more to do. That’s why he will keep fighting to lower drug costs, grocery prices, and energy bills.” Critical Perspective However, many Americans find these assurances lacking. The everyday experience at grocery stores and gas stations starkly contrasts with the administration’s optimistic declarations, leading to a disconnect between government rhetoric and public reality. As a Constitutional Republic dedicated to ensuring democracy and the well-being of its citizens, it is imperative that our government implements policies that stabilize the economy and reduce the financial burden on American families. The rising costs of essential goods and services are a pressing concern that requires immediate and effective action to safeguard the economic future of our nation. [Source: Poll, Whitehouse on X, BLS] https://d3nkl3psvxxpe9.cloudfront.net/documents/econTabReport_maqVHQt.pdf https://x.com/WhiteHouse/status/1800957041390792843 https://www.bls.gov/news.release/pdf/cpi.pdf #GoRightNews Shared by Peter Boykin - American Political Commentator / Citizen Journalist / Activist / Constitutionalist for Liberty Web: https://PeterBoykin.com Kick: http://Kick.com/PeterBoykin YouTube: https://youtube.com/PeterBoykinForAmerica Twitter: https://twitter.com/GoRightNews Telegram: http://t.me/realpeterboykin Rumble: http://Rumble.com/GoRightNews Like the Content? Please Support! - Go Right News: Stripe: https://gorightnews.com/donations/support-gorightnews/ Cash App: http://cash.app/\$PeterBoykin1”
Table D.9. Top-2 posts by total engagement for coordinated pages in Facebook
@@
”!You are a Golden Child! A Golden Child 94 Unicorn 94 Third Eye 94 Covenant 94 Harmony 94 Praise God 94 Divine Gene 94 Nine Six 94 Blue Eyes 94 Indigo Child 94 Ultra Maga 94 John John 94 White Hat 94 American Eagle 94 Carry On 94 New Earth 94 Pineapple 94 PassionForFruit”
@@
”#CAWildfireSituationUpdate as of 6-16-2024, 5:03 P.M. PST. #PostFire - 12,266 acres, 2% containment #HesperiaFire - 1,330, 7% #JunesFire - 1,076, 70% #JacksonFire - 876, ?#HernandezFire - 600, 25% #MaxFire - 500, 0% #LisaFire - 350, 0% #PointFire 150, 0%”
@@
”What is it like to be wiser than your Creator? Job tried that and later repented in dust and ashes. Job was a wise and blessed man of faith.”
@@
”What was the true history behind all of these melted ruins? Certainly not the result of erosion, but perhaps a plasma storm? The results of a sudden flip in our polar magnetic electric field perhaps.”
@@
”WAVELAND, Mississippi A boil advisory was issued Mon afternoon dt burst in the main water line. NASHVILLE, Indiana A boil advisory was issued Mon after a water main break on Honeysuckle Ln. MONTICELLO, Kentucky A boil advisory was issued Mon for the East Hwy 92 area dt a water main break. MILAN, Ohio Seminary Rd bw Perrin Rd/Broad St in Milan is closed dt a water main break. SASKATOON, SK - Canada North Park Wilson School is closed dt a water main break. NEW CASTLE CO., Delaware Shipley Rd at Foulk Rd was closed Mon night dt a water main break. MONTGOMERY CO., Maryland Dt a water main break Mon, a portion of the NW Branch Stream Valley Park near Highwood Terrace was undergoing repairs. SIOUX FALLS, South Dakota A water main break Mon afternoon caused flooding in downtown. BOURBON CO., Kansas There is a water main break at the Bourbon Co Transfer Station. FORTUNA, California Water will be shut off Wed to repair a broken water main on S Fortuna Blvd.”
@@
Table D.10. Top-5 posts by total engagement for cross-platform coordinated accounts
Appendix E Topic Analysis
We report the inferred topics and representative posts in each platform.
@@ Appendix E Topic Analysis
rfk ”If Kennedy iselected it will cost the evil doers over a trillion dollars in the first year. If Kennedy is in the debate it will show Biden as sinile, Trump as weak and Bobby will be elected. So they won’t let Bobby in but will compensate CNN and anyone that might go to jail.” ”I’ve never been more convinced that both Trump and Biden fear Kennedy. The uniparty has no soul, and they hate democracy. We will never forget their attempts to silence us. Kennedy FTW”
@@
debate ”Biden’s really not up to debate is what this means. Whenever a question is asked, Biden’s promoters will either tell the answer in an earplug or use a prompter screen on the podium so Biden can read the answer. This is funny.” ”Trump has guts though. They tried to make it as unfriendly for him as possible, and he still is going to do it. I’m sure they were hoping he would say no way, and then Biden could brag about Trump being afraid to debate him.”
@@
covid-vaccine ”Remember this on election day, remember All the ”suddenly died” family and friends! Biden did this, he made vaccines mandatory! Just one of many bad Biden decisions!” ”Trump was during Pandemic; FAILED to protect the American people when he knew damn well how infectious; deadly Covid Virus was, instead Trump told America ”Covid was a Hoax”; failed to create a National Suppression Plan to save lives 700KAmericans DIED under Trump!”
@@
bowman-latimer ”Bowman called Biden a liar. Bowman voted against the Infrastructure Act. Bowman even voted against the Debt Limit increase that could have jeopardized Social Security. Bowman did not earn another term as a Democrat.” ”Bowman is his worst enemy. His and Rape denying has turned his district against him. His Votes AGAINST the Biden Administration’s Progressive policies are NOT helping his district! Vote !”
@@
border-security ”with your lies. It is that BLOCKED Bipartisan Border Security Bill bc told GOP to ”Kill The Bill”. publicly stated it. The Border Crisis is NOW all on for Blocking Border Bill to SECURE THE BORDER!” ”This bill was a sham and would not close border,but would allow more to take American jobs!! Donalds reiterates why GOP rejected ‘bipartisan’ border bill to head off potential debate talking point”
@@
epstein-files ”His flights with Epstein were with family present and years before he had an island. He’s also the only one saying anything about releasing the logs, Biden sure isn’t, and Trump took umbrage when DeSantis started advocating for their release for some strange reason…” ” just another day in ’Murica, home of the child sex slaves. Trump made these folks feel they deserved Matt Gaetz’ job, Trumps’ job or Epstein’s job if they just ’networked’ enough. Investigate the real ’satanic’ sex cult that police and politicians hide.”
@@
student-loan ”This administration and president Biden seem to be answerable to nobody between giving money to student loan forgiveness even after the court ruled against them not getting Kennedy secret service protectiin, the list goes on no accountability.” ””Summary: We estimate that President Bidenś recently announced ”New Plans” to provide relief to student borrowers will cost \$84 billion, in addition to the \$475 billion that we previously estimated for President Bidenś SAVE plan.””
@@
hunter-biden ”Hunter Biden’s laptop has nothing to do with the false Electoral College votes sent in for the 2020 election. Also, Alexander Smirnov (an FBI informant) was INDICTED for his false accusations about Barisma and the Bidens.” ”Actions speak louder than how they phrased something – especially when the actions are repeated over and over again. Their lies are proven in this week’s Hunter Biden trial where the laptop is used as evidence.”
@@
blacks-against-biden ”People aren’t paying attention and don’t expect him to be the President and don’t know about Biden’s accomplishments. That’s what the polls show now but that’s gonna change. Trump isn’t gonna win 20% of the black vote. No were here near it” ”’At this stage in his presidency (3.5 years in) Biden must stop trying to convince Black voters his policies have made the US a significantly better place. Instead he should spend everything he’s got highlighting how anti-Black Trump is and how quickly any past gains will be lost.”
Table E.11. Topics and their representative posts on $\mathbb{X}$
@@
vaccine-skepticism ”BREAKING: US doctors and scientists are currently investigating whether the COVID-19 virus is to blame for an “unusual pattern” of rare and deadly cancers that have been popping up in the wake of the pandemic.GeneralMCNews” ”You, the unvaccinated are a special group of people. You stood your ground. You did not let the pressures of society sway you. You withstood the harshest discrimination seen in modern times. You were excluded from society. Some of you lost friendships and relationships with family members. You are the best that society has to offer. You used your analytical judgment and common sense. You are the tree that withstood the hurricane. You are superheroes! CovidVaccineTruth CovidVaccineTruthChat”
@@
flat-earth-conspiracy ”The Encyclopedia of Freemasonry reveals that the Earth is indeed flat. Organizations like The Freemasons know the truth about God and our flat earth but they choose to keep it hidden.” ”Flat Earth was the cosmology of our ancestors who had direct contact with the Divine because they were closer to the Divine than we are now. That is how the ancient’s knew things despite not having access to technology we have now.”
@@
Table E.12. Topics and their representative posts on Telegram
Topic Representative Posts
late-show
against-trump ”Stephen Colbert Speaks Truth to Trump’s Most Damaging Lie - #stephencolbert #TheLateShow #Trump #DonaldTrump #MAGA #immigrants” ”Stephen Colbert Speaks Truth to Trump’s Most Damaging Lie - Rick Strom #stephencolbert #TheLateShow #Trump #DonaldTrump #MAGA #immigrants”
barron-trump-rnc ”Former President Donald Trump said his youngest son Barron Trump was 17-years-old in an interview, when in fact he is 18 and a nominate delegate for the RNC.” ”Former President Donald Trump’s 18-year-old son, Barron Trump, is serving as a delegate to the Republican National Convention.”
2020-election-acceptance ”Republican National Committee co-chair Lara Trump said it’s ’obvious’ that Donald Trump has accepted the 2020 election results, when in fact, he has not.” ”Republican National Committee co-chair Lara Trump said it’s ’obvious’ that Donald Trump has accepted the 2020 election results, when in fact, he has not.”
trump-trial-dayoff ”Judge Juan Merchan gave Donald Trump the day off from his hush money trial to attend Barron Trump’s graduation ceremony, but evidence suggests that he is attending political fundraiser.” ”Judge Juan Merchan gave Donald Trump the day off from his hush money trial to attend Barron Trump’s graduation ceremony, but evidence suggests that he is attending political fundraiser.”
jacked-up-biden ”Rep. Greg Murphy, a Republican, says he has evidence that President Joe Biden was using ”something” before his popular State of the Union address.” ”Rep. Greg Murphy, a Republican, says he has evidence that President Joe Biden was using ”something” before his popular State of the Union address.”
trump-classified-documents ”More bombshell evidence dropped against former President Donald Trump in his classified documents case, just as Judge Aileen Cannon delayed the trial indefinitely.” ”More bombshell evidence dropped against former President Donald Trump in his classified documents case, just as Judge Aileen Cannon delayed the trial indefinitely.”
rfk ”Exciting Announcement! Join us for a special interview with independent presidential candidate RFK Jr. on Monday, June 17, immediately following The Young Turks! Tune in on tyt.com/live YouTube, Facebook, or Twitch! TYT members get exclusive access to a members-only Bonus Episode after the interview, where Cenk will ask even more personal questions and dive deeper into the issues! Sign up as a member today at to join the discussion! Don’t miss this unique opportunity to engage with crucial political dialogue and watch TYT hold a presidential candidate accountable! See you there!” ”Exciting Announcement! Join us for a special interview with independent presidential candidate RFK Jr. on Monday, June 17, immediately following The Young Turks! Tune in on tyt.com/live YouTube, Facebook, or Twitch! TYT members get exclusive access to a members-only Bonus Episode after the interview, where Cenk will ask even more personal questions and dive deeper into the issues! Sign up as a member today at to join the discussion! Don’t miss this unique opportunity to engage with crucial political dialogue and watch TYT hold a presidential candidate accountable! See you there!”
coulter-ramaswamy ”Conservative Ann Coulter told former Republican presidential candidate Vivek Ramaswamy said while she agrees with him on many issues, she wouldn’t vote for him because he’s “an Indian.”” ”Conservative Ann Coulter told former Republican presidential candidate Vivek Ramaswamy said while she agrees with him on many issues, she wouldn’t vote for him because he’s “an Indian.””
trump-falsifying-records ”Fox News’ Jeanine Pirro and conservative commentator Charlie Kirk are melting down over former President Donald Trump’s 34-count conviction for falsifying business records in his hush money trial.”
bannon-sentence ”Donald Trump adviser Steve Bannon proclaimed that ”́no one” will shut him up as he was ordered to report to prison for refusing to testify before congress.”
Table E.13. Topics and their representative posts on Facebook
Appendix F Characterization of AI Generated Content
Regarding the AI-Generated Content (AIGC) results, we conducted a qualitative analysis to characterize AIGC across platforms. Specifically, we analyzed the top 20 AIGC posts by popularity from each platform and used GPT-4 with zero-shot learning to estimate a partisan score. Our findings show that Telegram has an average partisan score close to zero, while $\mathbb{X}$ and Facebook exhibit higher scores of 0.5 and 0.4, respectively. This highlights differences among AIGC diffusion between the three platforms that help to interpret and characterize the results.